# Safe In-Context Reinforcement Learning

### v1

@@ 3 Safe In-Context Reinforcement Learning
$\textstyle\pi_{t+1}\in\operatorname*{argmax}_{\pi}L_{\Sigma}(\pi,\lambda_{t})\,\quad\lambda_{t+1}=[\lambda_{t}+\eta\,\max_{k}g_{k}(\pi_{t+1})]_{+}\quad\eta>0.$ (7)
@@
The expected return $\mathbb{E}_{\pi}\quantity[\sum_{k=1}^{K}G(\tau_{k})]$} and expected cost $\mathbb{E}_{\pi}\quantity[G_{c}(\tau_{k})]$} are bounded and continuous in $\pi$. The constrained problem (4) admits an optimal feasible policy $\pi^{\star}$.
@@
[Paternain (2018)] Problem (4) satisfies conditions ensuring zero duality gap with (6) and there exists optimal Lagrange multipliers ${\bm{\lambda}}^{\star}\succeq 0$ such that
$\textstyle\underset{\text{s.t.}\ \forall k,{\mathbb{E}_{\pi}\quantity[G_{c}(\tau_{k})]}\leq\delta}{\max_{\pi}}{\mathbb{E}_{\pi}\quantity[\sum_{k=1}^{K}G(\tau_{k})]}\ \;=\;\min_{{\bm{\lambda}}\succeq 0}\max_{\pi}L(\pi,{\bm{\lambda}})$
We now show that the set of the fixed points of (7) and the set of the optimizers of (4) are equal.
@@
[Proof in Appendix B.1] We say a pair $(\bar{\pi},\bar{\lambda})$ is a fixed point of (7) if $\bar{\pi}\in\arg\max_{\pi}L_{\Sigma}(\pi,\bar{\lambda})$, and for all sufficiently small $\eta>0$, $\bar{\lambda}=[\bar{\lambda}+\eta\,\max_{k}g_{k}(\bar{\pi})]_{+}$. Let Assumptions 1 - 2 hold. Then every fixed point $(\bar{\pi},\bar{\lambda})$ of (7) is feasible and primal-optimal for (4). Conversely, every primal-optimal policy $\pi^{*}$ admits a corresponding fixed point $(\pi^{*},\bar{\lambda})$ for any multiplier $\bar{\lambda}\geq\|{\bm{\lambda}}^{\star}\|_{\infty}$, where $\bm{\lambda}$$^{\star}$ is an optimal dual solution.
To implement the policy update (7), we consider an actor-critic approach following Grigsby et al. (2024a). The pseudocode is provided in Algorithm 1. Since our update (7) resembles an exact penalty convex optimization technique when $\lambda\geq\|{\bm{\lambda}}^{\star}\|_{\infty}$ (Nocedal & Wright, 2006), we call the resulting actor-critic algorithm Exact Penalty Policy Optimization (EPPO).
@@ 4 Experiments
We now empirically investigate the proposed safe ICRL algorithms. Specifically, we aim to answer the following two questions.
@@
    Do the proposed safe ICRL algorithms demonstrate generalization to out-of-distribution (OOD) tasks in complex environments with safety constraints?
@@
    How effectively do the proposed safe in-context RL algorithms achieve flexible reward-cost tradeoffs under varying cost tolerance?
We evaluate our approach on two constrained environments: (1) SafeDarkRoom and (2) SafeDarkMujoco: Point and Car.
@@
These are modified from DarkRoom (Laskin et al., 2023) and SafetyGym benchmark (Ji et al., 2023) respectively. SafeDarkRoom is a grid-world setting where the agent can only perceive its own position and lacks visibility of the goal or obstacle locations. Rewards are sparse, with a positive reward obtained upon reaching the single goal in the map. Costs arise from multiple obstacles scattered across the environment. For instance, 25 obstacles in a $\times9$ grid map, which incur a cost of 1 each time the agent steps on one of the obstacles. To succeed, the agent must learn obstacle positions based on encountered cost signals, introducing additional complexity compared to goal-only environments due to the presence of multiple hazards. SafeDarkMujoco environment operates in continuous space with MuJoCo simulation (Todorov et al., 2012) but with a setup analogous to SafeDarkRoom. The robot senses its internal physical states, such as velocity, acceleration, position, and rotation angle, while all lidar inputs are turned off, preventing detection of obstacles and the goal. Consequently, the agent must develop exploratory behaviors to identify obstacles and goals, akin to SafeDarkRoom design, while navigating to the goal and minimizing collisions by learning from the previous collisions in the context.
Measuring OOD generalization.
@@
To measure the OOD generalization of our proposed safe ICRL methods, we employ a challenging distance-based obstacle and goal spawning strategy: center-oriented for pretraining and edge-oriented for evaluation. The agent spawns at the map center, with obstacles and the goal distributed proportionally closer to this center during pretraining. During evaluation, obstacles and the goal follow an edge-oriented distribution. This approach applies similarly to the SafeDarkRoom and SafeNavigation environments. We argue that this setup is more challenging than those in prior studies (Laskin et al., 2023; Zisman et al., 2023; Son et al., 2025), as the agent must learn extrapolation rather than interpolation. More precisely, during pretraining, the grid position $(i,j)$ has an obstacle with probability $\mathbb{P}_{\text{train}}((i,j))\propto e^{-\alpha d((i,j),c)}$, where $c$ is the map center, $d(\cdot,\cdot)$ denotes Euclidean distance, and $\alpha>0$ to promote central density.
@@
We now demonstrate emergent safe learning behaviors in OOD test tasks, thus giving an affirmative answer to Question (i). We use supervised pretraining and reinforcement pretraining in Section 3 on pretraining tasks with center-oriented obstacles and goals. Particularly, for supervised pretraining, the source algorithm to generate the dataset is PPO-Lagragian (Ray et al., 2019). goals. For reinforcement pretraining, we update the model for 30,000 steps on SafeDarkRoom and 10,000 steps on SafeDarkMujoco using Algorithm 1.
After pretraining, we evaluate the learned agents in evaluation tasks with edge-oriented obstacles and goals. The agent is able to interact with the evaluation task for multiple episodes, but cannot update its network parameters. During evaluation, for supervised pretraining, we set RTG to 1.0 and CTG to 0.0, targeting successful performance while minimizing cost violations. Reinforcement pretraining is solely conditioned on CTG. To evaluate robustness to CTG values during evaluation, we set CTG to a value uniformly sampled from the interval $[1,15]$ for SafeDarkRoom and $[10,50]$ for SafeDarkMujoco. Precisely, we sample $100$ distinct CTG values and report the average performance over these $100$ distinct CTG values. We recall that although the safe supervised pretraining method we propose is novel, we mostly regard it as a straightforward baseline. As can be seen in Figure 1, safe supervised pretraining succeeds in SafeDarkRoom as the episode return grows and the episode cost decreases. This aligns with the previously observed success of algorithm distillation in DarkRoom (Laskin et al., 2023; Zisman et al., 2023). But this baseline approach fails in SafeDarkMujoco. We conjecture that this is because of insufficiency in the dataset, as supported by ablation studies (Figure 6). By contrast, our safe reinforcement pretraining approach consistently succeeds in all three tested domains. It is also worth highlighting that in SafeDarkRoom, the safe reinforcement pretraining enables much faster adaptation to the new evaluation task than the safe supervised pretraining approach.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption(a) Reinforcement Pretraining(b) Supervised PretrainingSafeDarkroomSafeDarkMujoco-PointSafeDarkMujoco-Car
Figure 1: Evaluation performance of superivsed and reinforcement pretraining. The curves are averaged over 1616 distinct OOD evaluation tasks, with each task featuring edge-oriented goal and obstacles. The shaded regions indicate standard errors. The $x$-axis is the episode index $k$ during the evaluation task. The $y$-axis is the corresponding episode return $G(\tau_{k})$ and the epsiode cost $G_{c}(\tau_{k})$ respectively. The straightforward safe supervised pretraining baseline succeeds only in SafeDarkRoom, while our novel safe reinforcement pretraining method succeeds in all three domains.
@@
We now demonstrate that we can adjust the behavior of the pretrained policy in test time by simply changing the RTG and/or CTG without making any parameter updates. Intuitively, with a higher CTG (i.e., a higher cost tolerance), the agent should afford bolder exploration and thus obtain higher rewards. Conversely, a lower CTG should promote conservative policies that prioritize safety, potentially at the expense of suboptimal rewards.
For the supervised pretrained model, we initialize RTG at 0.5 and CTG at 0.0, then adjust RTG = $\max\{0.5,\frac{\text{CTG}}{10}\}$ as CTG increases from 5 to 15 for SafeDarkRoom and 0 to 10 for SafeDarkMujoco. This setup enables gradual increases in both RTG and CTG to demonstrate balanced reward-cost trade-offs, consistently applied to SafeDarkRoom and SafeDarkMujoco. However, the supervised pretrained model fails to learn trade-off behaviors, exhibiting random patterns with varying RTG and CTG (Figure 2.(b)). We hypothesize that this stems from insufficient CTG diversity during training, leading to dataset inadequacy.
In our reinforcement pretrained model, we focus on controlling the CTG exclusively. This design simplifies the regulation of RTG based on the CTG. Simultaneously controlling both RTG and CTG is challenging, as determining the feasibility of RTG/CTG pairs is complex (Liu et al., 2023). Instead, the reinforcement pretrained model is designed to control CTG and pursue the maximal possible return for a given CTG, eliminating the need to justify the feasibility of specific solutions. Figure 2.(a) illustrates this relationship, showing that as cost limits increase, the return also increases.
Supervised pretraining depends on a long replay buffer to learn from logged learning histories. Due to context length constraints, we need to subsample trajectories from these learning histories for supervised pretraining. These histories are sampled from distinct learning phases: early (low return, high cost), middle (average return, average cost), and expert (high return, low cost). To make sequence modeling feasible, trajectories need to be labeled with RTG, which provides information about the expected phase (Dai et al., 2024). However, this approach requires significantly more data, as noted by Liu et al. (2023). In contrast, reinforcement pretraining leverages online interaction histories, eliminating the need to distinguish between different data phases.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption(a) Reinforcement Pretraining(b) Supervised PretrainingSafeDarkroomSafeDarkMujoco-PointSafeDarkMujoco-Car
Figure 2: Evaluation performance of supervised and reinforcement pretraining with varying CTG. The curves are displayed across a range of cost limits. For each cost limit, the result is averaged over 1616 distinct OOD evaluation tasks, with each task incorporating edge-oriented goals and obstacles. The shaded regions indicate standard errors. The $x$-axis is the CTG. The $y$-axis is the total episode return (i.e., $\sum_{k=1}^{K}G(\tau_{k})$) and the maximum episode cost (i.e., $\max_{k\in K}G_{c}(\tau_{k})$) with the corresponding CTG. The policy from safe reinforcement pretraining succeeds in converting a higher cost limit (i.e., a higher CTG) to a higher return while the policy from safe supervised pretraining fails to do so.
Ablations.
We further conduct ablation studies on reinforcement pretraining and supervised pretraining, examining shared factors such as context length and model size, alongside specific factors such as the dataset size for supervised pretraining. Our findings show distinct sensitivities to these factors. Reinforcement pretraining is largely unaffected by model size, whereas model size significantly influences supervised pretraining performance. For context length, reinforcement pretraining performs better with longer sequences, while supervised pretraining performs worse. Conversely, supervised pretraining excels with shorter context lengths, while reinforcement pretraining performs poorly. These results suggest that learning long-term credit assignment is more challenging in offline reinforcement learning due to dataset constraints, whereas online learning, with access to environment interactions, handles long-term credit assignment more effectively. We confirm the well-established finding that supervised pretraining is highly sensitive to dataset size. When the dataset is small, the model fails to learn, exhibiting random behavior on out-of-distribution data. In Figure 5.(c), we compare EPPO to a naive primal–dual solver for (6) with per-episode multipliers $\lambda$. The overall structure of this naive algorithm resembles RCPO (Tessler et al., 2018). EPPO trains more stably with minimal tuning and solves SafeDarkRoom and SafeDarkMujoco without changes beyond time-limit related settings (Table 1). The detailed evaluation curves of the ablation study are provided in Figures 56 in Appendix E. We evaluate a variant of safe supervised pretraining incorporating algorithm distillation with noise in Appendix D. The results indicate that learning histories generated by noisy actions do not effectively represent true learning histories in complex environments with safety constraints.

### v2

@@ 4 SCARED: Safe Contextual Adaptive Reinforcement Pretraining
$\textstyle\pi_{t+1}\in\operatorname*{argmax}_{\pi}L_{\Sigma}(\pi,\lambda_{t})$ (6)
$\lambda_{t+1}=[\lambda_{t}+\eta\,\max_{k}g_{k}(\pi_{t+1})]_{+}\quad\eta>0.$
We analyze the updates in (6) under the following regularity assumptions.
@@
The expected return $\mathbb{E}_{\pi}\hskip-1.0pt\quantity[\sum_{k=1}^{K}\hskip-1.0ptG(\tau_{k})]$} and expected cost $\mathbb{E}_{\pi}\hskip-1.0pt\quantity[G_{c}(\tau_{k})]$} are bounded and continuous in $\pi$. The constrained problem (4) admits an optimal feasible policy $\pi^{\star}$.
@@
[Paternain (2018)] Problem (4) satisfies conditions ensuring zero duality gap with (4) and there exist optimal Lagrange multipliers ${\bm{\lambda}}^{\star}\succeq 0$ such that
$\underset{\text{s.t.}\ \forall k,{\mathbb{E}_{\pi}\hskip-1.0pt\quantity[G_{c}(\tau_{k})]}\leq\delta}{\max_{\pi}}{\mathbb{E}_{\pi}\hskip-1.0pt\quantity[\sum_{k=1}^{K}\hskip-1.0ptG(\tau_{k})]}\ =\min_{{\bm{\lambda}}\succeq 0}\max_{\pi}L(\pi,{\bm{\lambda}})$
The following theorem shows that the set of the fixed points of (6) and the set of the optimizers of (4) are equal, meaning that the objective of our approach aligns with the safe ICRL problem statement.
@@
[Proof in Appendix B.1] We say a pair $(\bar{\pi},\bar{\lambda})$ is a fixed point of (6) if $\bar{\pi}\in\arg\max_{\pi}L_{\Sigma}(\pi,\bar{\lambda})$, and for all sufficiently small $\eta>0$, $\bar{\lambda}=[\bar{\lambda}+\eta\,\max_{k}g_{k}(\bar{\pi})]_{+}$. Let Assumptions 1 - 2 hold. Then every primal-optimal policy $\pi^{*}$ admits a corresponding fixed point $(\pi^{*},\bar{\lambda})$ for any multiplier $\bar{\lambda}\geq\|{\bm{\lambda}}^{\star}\|_{\infty}$, where $\bm{\lambda}$$^{\star}$ is an optimal dual solution. Conversely, every fixed point $(\bar{\pi},\bar{\lambda})$ of (6) is feasible and primal-optimal for (4).
Our update (6) resembles an exact penalty convex optimization technique when $\lambda\geq\|{\bm{\lambda}}^{\star}\|_{\infty}$ (Nocedal & Wright, 2006), hence the naming. We implement it with an actor-critic approach following Grigsby et al. (2024a); see Algorithm 1 for pseudocode. Separate critics estimate the reward Q-function $Q_{\theta_{v}}$ and cost Q-function $Q^{c}_{\theta_{c}}$, trained with TD targets from target networks. The actor maximizes $Q_{\theta_{v}}$ while penalizing $Q^{c}_{\theta_{c}}$ for episodes exceeding the budget.
@@ 5 Experiments
We now empirically investigate the proposed safe ICRL algorithms, aiming to answer the following research questions:
@@
    Does SCARED generalize to both out-of-distribution tasks and unseen in-distribution (ID) tasks in safety-constrained environments?
@@
    How effectively does SCARED achieve flexible reward-cost tradeoffs under varying safety budgets at test time, without parameter updates?
5.1 Baselines
We consider two baseline classes: algorithm distillation for safe RL, which encodes safety into the policy’s forward pass without test-time parameter updates, and safe meta-RL methods, which adapt via parameter updates at evaluation but do not use in-context interaction histories for action selection.
Algorithm Distillation for Safe RL. We adopt algorithm distillation for safe RL (safe AD) as a strong baseline for safe ICRL. Following algorithm distillation (Laskin et al., 2023), we collect a dataset $\{\Xi_{i}\}$ by executing existing safe RL algorithms (e.g., (Tessler et al., 2019; Ray et al., 2019)) across safety-constrained environments (e.g., (Ji et al., 2023; Gu et al., 2025)). As a result, each trajectory $\Xi_{i}$ consists of multiple episodes and each episode contains both the reward and the cost (cf. (3)).
To ensure a fair and competitive baseline, the policy is conditioned on the full in-context interaction history $H^{k}_{t}$, which includes per-step reward and cost signals, as well as on CTG and return-to-go (RTG), a target specifying the desired future reward that steers the policy toward a desired reward level (Liu et al., 2023). The raw reward and cost information in $H^{k}_{t}$ enables algorithm distillation (Laskin et al., 2023), allowing the model to imitate adaptation behavior present in the offline trajectories. In contrast, RTG and CTG serve a distinct role as conditioning variables that bias action selection toward desired reward-cost tradeoffs at test time. Formally, given a trajectory $\Xi=(\tau_{1},\dots,\tau_{K})$, the safe AD objective at state $S_{t}^{k}$ is $\log\pi_{\theta}\!\left(A_{t}^{k}\mid S_{t}^{k},H_{t}^{k},G_{t}(\tau_{k}),G_{c,t}(\tau_{k})\right)$. During pretraining, RTG and CTG are available since episodes are complete. At test time, they are replaced by user-specified targets, allowing control over the reward-cost tradeoffs without parameter updates. Safe AD relies on a large replay buffer of logged learning histories, which must be subsampled due to context-length constraints and labeled with RTG to distinguish between early, intermediate, and expert phases of training (Dai et al., 2024). We also evaluate a noise-based variant of Safe AD (Zisman et al., 2023); however, training on optimal trajectories with noise fails to provide the behavioral diversity required for OOD adaptation, highlighting the learning mechanism in our Safe AD baseline for reliable in-context safety adaptation (Appendix D).
Safe Meta RL. We compare our approach against gradient-based safe meta RL methods, specifically MAML with penalty (Finn et al., 2017), and SafeMeta (Xu & Zhu, 2025). MAML with penalty optimizes a model-agnostic initialization via cost-weighted penalties that can be adapted to new tasks with a few gradient steps. SafeMeta combines a closed-form, one-step safe policy adaptation method with a Hessian-free meta-training algorithm to guarantee anytime safety. By design, these methods adapt through per-task parameter updates rather than conditioning on cross-episode interaction histories.
5.2 Experimental Setup
Safe ICRL Benchmarks.
We evaluate our approach on three constrained environments, categorized into two complementary generalization regimes: (i) structural OOD tasks and (ii) unseen ID tasks. First, we use SafeDarkRoom and SafeDarkMujoco (Point and Car) to evaluate adaptation to OOD tasks with shifted goal and obstacle configurations. Second, we evaluate on SafeVelocity (HalfCheetah and Ant), representing the unseen ID task regime commonly used in meta-RL benchmarks. We discuss the difficulty gap between these two categories in Section 5.3.
@@ 5.2 Experimental Setup
The first two environments are modified from DarkRoom (Laskin et al., 2023) and SafetyGym benchmark (Ji et al., 2023) respectively. SafeDarkRoom is a grid-world setting where the agent can only perceive its own position and lacks visibility of the goal or obstacle locations. Rewards are sparse, with a positive reward obtained upon reaching the goal in the map. Costs arise from multiple obstacles scattered across the environment. For instance, 25 obstacles in a $\times9$ grid map, which incur a cost of 1 each time the agent steps on one of the obstacles. To succeed, the agent must learn obstacle positions based on encountered cost signals, introducing additional complexity compared to goal-only environments due to the presence of multiple hazards.
SafeDarkMujoco operates in continuous space with MuJoCo simulation (Todorov et al., 2012), with a setup analogous to SafeDarkRoom. The robot senses its internal physical states, such as velocity, acceleration, position, and rotation angle, while range-based sensing (e.g., lidar) is unavailable, preventing direct detection of obstacles and the goal. Consequently, the agent must develop exploratory behaviors to identify obstacles and goals, akin to the SafeDarkRoom design, while navigating to the goal and minimizing collisions by learning from previous collisions in the context. SafeVelocity follows the design of Ji et al. (2023); however, we vary the target velocities and hold out a set of unseen velocities for evaluation. Such velocity-conditioned setups are often considered within meta-RL frameworks (Finn et al., 2017), where different target velocities are sampled from a fixed range (e.g., $[0,1]$). We thus consider SafeVelocity tasks as unseen ID tasks, which are more straightforward than OOD generalization.
Training Setup. For OOD task environments, all algorithms are trained on source tasks with center-oriented obstacles and goals. For unseen ID task environments, we hold out a set of target velocities multipliers uniformly sampled from $[0.5,1.0]$ in increments of $0.05$ during training and evaluating performance on these held-out velocities at test time.
Safe AD is trained on trajectories generated using PPO-Lagrangian (Ray et al., 2019). For SCARED, we perform online reinforcement pretraining for 30,000 steps on SafeDarkRoom and 10,000 steps on SafeDarkMujoco using Algorithm 1, and we train for 1,000 steps on the SafeVelocity environments (HalfCheetah and Ant). Safe meta-RL baselines are trained for up to 15,000 steps on each environment, with most converging earlier, using their respective update rules. For safe AD and safe meta-RL baselines, we perform hyperparameter optimization to ensure fair comparison across methods. Additional implementation details and used hyperparameters are provided in Appendix C.
Evaluation Setup. To answer the aforementioned research questions, we use a consistent setup for CTG and RTG during evaluation. For SCARED, we evaluate robustness to safety budgets by conditioning on CTG values sampled uniformly from $[1,10]$ in SafeDarkRoom, $[10,50]$ in SafeDarkMujoco, and $[0,5]$ in SafeVelocity, reporting performance averaged over 100 distinct CTG targets. This validates that pretraining with diverse CTG targets enables budget-conditioned generalization to OOD tasks. The same intervals and sample counts are used for the cost budgets of SafeMeta and MAML with penalty. For Safe AD, we condition the policy on paired target values. Specifically, we initialize RTG at $0.1$ and CTG at $0.0$, and then adjust RTG according to $\mathrm{RTG}=\max\left\{0.1,\frac{\mathrm{CTG}}{\mathrm{CTG}_{\max}}\right\}$ as CTG increases from 0 to $\mathrm{CTG}_{\max}$, where $\mathrm{CTG}_{\max}=10$ for SafeDarkRoom, $50$ for SafeDarkMujoco, and $5$ for SafetyVelocity. Performance is reported as an average over 100 paired RTG-CTG targets.
5.3 Measuring OOD generalization
@@
To measure the OOD generalization of our proposed safe ICRL methods, we employ a challenging distance-based obstacle and goal spawning strategy: center-oriented for pretraining and edge-oriented for evaluation. The agent spawns at the map center, with obstacles and the goal distributed proportionally closer to this center during pretraining. During evaluation, obstacles and goals follow an edge-oriented distribution. This approach applies similarly to the SafeDarkRoom and SafeDarkMujoco environments. We argue that this setup is more challenging than those in prior studies (Laskin et al., 2023; Zisman et al., 2023; Son et al., 2025), as the agent must learn extrapolation rather than interpolation. More precisely, during pretraining, the grid position $(i,j)$ has an obstacle with probability $\mathbb{P}_{\text{train}}((i,j))\propto e^{-\alpha d((i,j),c)}$, where $c$ is the map center, $d(\cdot,\cdot)$ denotes Euclidean distance, and $\alpha>0$ to promote central density.
@@
Proposition 1 shows that as $\alpha\to\infty$, the overlap between the training and test distributions vanishes. As a result, the model must generalize to inputs far outside the regions observed during training.
5.4 Analysis
@@
Figure 1: Evaluation performance under increasing in-context episodes. Curves are averaged over environment-specific OOD evaluation tasks; shaded regions denote standard errors. The $x$-axis shows the episode index $k$, and the $y$-axis reports the episode return $G(\tau_{k})$ (top) and episode cost $G_{c}(\tau_{k})$ (bottom). SCARED shows consistent improvement in return and reduction in cost as the in-context interaction grows across all environments, and outperforms all baselines. Safe AD shows environment-dependent performance, failing in the SafeVelocity domain and performing competitively in others, but does not achieve the same performance as SCARED. Safe meta-RL algorithms fail to achieve safe learning behavior due to a lack of in-context interaction, despite updating parameters during evaluation.
We now demonstrate emergent safe learning behaviors in the test environments, thus giving an affirmative answer to Question (i). For OOD evaluation, we use edge-oriented obstacles and goals in SafeDarkRoom and SafeDarkMujoco. For unseen ID tasks, we evaluate on SafeVelocity (HalfCheetah and Ant), where target velocities are unseen during training. During evaluation, agents interact with each task for multiple episodes without updating parameters (except for meta-RL baselines, which adapt via parameter updates).
Figure 1 shows SCARED consistently outperforms all baselines by learning safe behavior through in-context interaction, showing increasing episode return and decreasing episode cost as the interaction history grows across all evaluated environments.
Safe AD performs competitively in SafeDarkRoom, showing a similar learning pattern but with slower and weaker adaptation than SCARED. However, its performance degrades as environment complexity increases. In SafetyAnt, episode cost rises while return declines through in-context episodes, and in SafeDarkMujoco, although return improves, cost does not decrease, failing safe adaptation.
Among safe RL baselines, MAML with penalty fails to achieve safe and high-performing adaptation across the evaluated environments. In contrast, SafeMeta achieves higher returns in SafetyVelocity domains but does not reduce safety violations, even when allowed to update parameters during evaluation. This suggests that parameter adaptation alone, without cross-episode interaction histories, is insufficient to regulate safety during online exploration.
@@
Figure 2: Evaluation performance under varying CTG targets. Results are averaged over environment-specific OOD evaluation tasks; shaded regions denote standard errors. The $x$-axis shows the target CTG, and the $y$-axis reports the total return $\sum_{k=1}^{K}G(\tau_{k})$ (top) and the maximum episode cost $\max_{k\in[1,K]}G_{c}(\tau_{k})$ (bottom). SCARED shows a clear pattern: as the cost budget increases, it achieves higher total return across all environments. Other algorithms fail to present this budget-adjustment behavior.
We demonstrate that the behavior of the pretrained policy can be adjusted at test time by varying CTG, without updating parameters. Intuitively, higher CTG values correspond to greater cost tolerance and enable more aggressive exploration and higher return. Conversely, lower CTG values induce more conservative, safety-focused behavior, potentially at the expense of suboptimal rewards. To study this, we use the same range of CTG introduced before.
SCARED controls CTG exclusively and pursues the maximal achievable return subject to the specified cost budget, avoiding the need to reason about the feasibility of arbitrary RTG-CTG pairs, which is known to be challenging in constrained settings (Liu et al., 2023). Figure 2 shows that increasing the cost budget consistently leads to higher return, indicating that SCARED learns to map the cost budget to an exploration strategy during in-context adaptation. Notably, maximum episode costs track the CTG target closely, with nearly all episodes respecting the specified budget.
In contrast, Safe AD fails to adapt consistently to varying CTG values, showing inconsistent behavior. For example, in SafeDarkRoom, Safe AD achieves relatively high total return at low CTG targets but shows a downward trend as the cost budget increases, failing to control return-cost tradeoffs. Similarly, safe meta-RL baselines do not show comparable budget-adjustment behavior. Regardless of specified CTG targets, their performance does not vary with the cost budgets, suggesting that parameter adaptation alone does not provide test-time control over reward-cost tradeoffs.
Ablations. We conduct ablations on SCARED and Safe AD over shared factors (context length, model size) and method-specific factors (safe AD dataset size), with results reported in Appendix E, revealing distinct sensitivities to these factors. SCARED is largely unaffected by model size, whereas model size significantly influences safe AD performance. For context length, SCARED performs better with longer sequences, while safe AD performs worse. Conversely, safe AD excels with shorter context lengths, while SCARED performs poorly. These results suggest that long-term credit assignment is harder in offline RL due to dataset constraints, while online learning benefits from direct environment interaction. We confirm that Safe AD is highly sensitive to dataset size: with limited data, it fails to learn and shows random behavior on OOD tasks.